OpenStudy (anonymous):

The cost of two different mobile phone contracts: Company A: $30 and $0.15 per minute for phone calls Company B: $10 and $0.55 per minute for phone calls When is company A cheaper than company B?

OpenStudy (anonymous):

holy crap I had this same problem in MTH 163. Comp. A 0.15x+30 Comp. B 0.55x + 10 0.15x+30=0.55x+10 30=0.40x+10 20=0.40x x=50 so at 50 minutes they are equal. lets test 45 minutes.. 0.15(45)+30=36.75 0.55(45)+10=34.75 at less than 50 minute company B is cheaper. so it has to be x>50 for company A to be cheaper.

OpenStudy (anonymous):

the function is A(t) = 30 + 0.15t B(t) = 10 + 0.55t t = minutes when A costs as much as B : A = B 30 + 0.15t = 10 + 0.55t 20 = 0.4t t = 50 then plug any number that's less than 50, let's plug 1 A(1) = 30 + 0.15 = 30.15 B(1) = 10 + 0.55 = 10.55 That means, when the t<50 , A costs more than B. Therefore A is cheaper then B when t>50

OpenStudy (anonymous):

prettier way of explaining it :P